The owner of Nuts2U Snack Shack mixes cashews worth $6.25 a pound with peanuts
worth $2.00 a pound to get a half-pound mixed nut bag worth $1.70. How much of each kind of nut is included in the mixed bag?

Answer by stanbon(75887)

The owner of Nuts2U Snack Shack mixes cashews worth $6.25 a pound with peanuts
worth $2.00 a pound to get a half-pound mixed nut bag worth $1.70.
How much of each kind of nut is included in the mixed bag?
---------------------------------
Note: Something is wrong with your data. You cannot mix
6.25 and 2.00 nuts and get 1.70 nuts.
---
Cheers,
Stan H.
--------
If you try to get a solution you end up with a negative quantity.
---
Value Equation: 6.25c + 2.00p = 1.70(1/2)
Quantity Equation: c + p = 1/2
------------------------------------
Multiply thru the 1st by 100
Multiply thru the 2nd by 200
-------------------------------
625c + 200p = 170*(1/2)
200c + 200p = 200*(1/2)
------
Subtract 2nd from 1st and solve for "c":
425c = -30(1/2)
c = -0.0353, which is unrealistic.
=====================================
Cheers,
Stan H.